# Radiation forces and torques in optics and acoustics

### v1

@@ III.3 Isotropic Rayleigh particles
Here the complex-valued quantities are the corresponding polarizablities of the particle, which are defined such that all of them have the dimension of volume: . Also, we wrote relations (III.3) using the general form with the polarizability matrix, which is diagonal in this case. Since there are different conventions for definitions of the electromagnetic and acoustic dipole and monopole moments, we present the fields produced by the moments adopted in our review in Appendix A.
The polarizablities can be connected with the scattering, absorption, and extinction cross sections, related as . In particular, for all polarizabilities in Eqs. (III.3), we have:
@@
These relations follow from the direct calculation of the ‘mixed’ and ‘scattered’ energy fluxes through a closed surface around the particle, normalized to the absolute value of the ‘incident’ energy flux density, assuming an incident plane-wave field [548, 320, 543, 107]. For particles without gain, and , where the equality is achieved in the lossless case. This imposes constraints on the polarizabilities, which are sometimes referred to as the optical theorem [407, 72, 437]. Note also that the absorption cross section is equal to the normalized absorption rate: , in the case of an incident plane wave.
@@
Explicit equations for the isotropic Rayleigh-particle polarizabilities in terms of the material parameters can be derived from the electromagnetic Mie theory and its acoustic analogue. The results are summarized in Appendix B and Table 4.
@@
Substituting Eqs. (III.3) into the ‘mixed’ terms (linear in the dipole and monopole moments) in Table 2, we obtain the absorption rates, forces, and torques on the particle conveniently expressed via the energy, canonical momentum, and spin densities (2), (4), (5), (7), (9), (10), (11) of the incident wave field:
@@
These are the most important optical and acoustic forces on small particles, which are schematically shown in Fig. 7. Remarkable symmetry between the optical and acoustic quantities justifies the unified framework we use. In Eqs. (22) one can distinguish gradient forces determined by gradients of the corresponding energy-density terms and radiation-pressure forces determined by the corresponding canonical momentum density terms. Note that in paraxial optical fields, the energy density gradients are directed near-orthogonally to the canonical momentum (propagation) direction, so that the gradient and radiation-pressure forces act in near-orthogonal directions. The torques (23) are naturally determined by the corresponding spin densities. Importantly, Eqs. (21)–(23) are valid for an arbitrarily inhomogeneous (structured) incident field, including, e.g., near-fields and evanescent waves.
@@
The forces and torques (22) and (23) underpin optical and acoustic manipulation of small particles, including trapping, tweezers, spanners, levitation, etc. [181, 387, 248, 109, 38, 579, 178, 523, 219, 607, 560, 234, 421, 428, 362, 610, 600, 117, 375, 412, 538, 62, 341, 211, 328, 184, 271, 585, 179, 318, 241, 121, 619, 489] Pioneering works on the dipole radiation forces were made by Gordon and Ashkin in 1980 [244, 44] (considering interaction of an electric-dipole particle with an incident electromagnetic wave) and by Gor’kov in 1961 for the acoustic gradient force [245]. Since then, there were numerous works extending these results [137, 402, 465, 181, 523, 70, 87, 100, 117, 543, 501, 488], until they acquired the modern form (22). Surprisingly, equations for optical and acoustic torques (23) were derived only recently [138, 126, 87, 100, 401, 501, 543].
@@
In a lowest-order approximation, the imaginary parts of the polarizabilities in Eqs. (21)–(23) are typically associated with the absorption in the particle. However, according to Eqs. (III.3), they are in fact determined by the extinction cross-sections: which include both absorption and scattering. This results in nonzero imaginary parts of the polarizabilities even for lossless particles with : , Eqs. (III.3) (see Appendix B). In particular, this means that the radiation-pressure forces in Eqs. (22) do not vanish in this lossless case [255]. This effect is sometimes called “radiation friction” [509, 402, 508, 12, 320, 543]. In contrast, the absorption rates and torques on lossless isotropic particles do vanish, but showing this requires taking into account the contributions from the ‘scattered’ fluxes, which are considered in the next Section III.4.
@@
It should also be noticed that a number of works on optical forces [13, 362, 288, 219, 565, 402, 523] use an alternative yet equivalent form of the electromagnetic radiation-pressure force in Eq. (22). It is obtained by expressing the canonical momentum densities via the difference of the kinetic Poynting momentum (3) and curls of the corresponding spin densities, according to the Belinfante-Rosenfeld relation (1): . However, such an artificial separation meant to introduce a term proportional to the Poynting vector and an additional “spin-curl” force can be misleading. Indeed, the canonical momentum density has a clear physical meaning as a natural measure of the phase gradient and the local wavevector (multiplied by the field intensity) [76, 78, 87, 100], universal for any wavefield [103], and independent of the polarization in paraxial uniformly-polarized fields. In contrast, the Poynting-vector and spin-curl contributions introduce two polarization-dependent force terms even in the simplest paraxial fields which partially cancel each other, resulting in the same radiation pressure force determined by the mean phase gradient in the field.
@@ III.4 Recoil force and torque
We now consider the second-order forces and torques, which appear from pure scattering-field fluxes and are quadratic in the dipole and monopole moments. The corresponding terms in Table 2 are denoted as “scattered (recoil)”. Figure 8 schematically explains the nature of the recoil force. It is related to the asymmetry of the scattered momentum flux, which originates from the interference of the symmetric electric and magnetic dipole or acoustic dipole and monopole contributions [402, 513, 567] (see also Appendix D).
@@
where and are the natural complexifications of the electromagnetic and acoustic kinetic momentum densities (3) and (8).
@@
The recoil force (24) has been examined in optics since 2009 [138, 402, 70, 87, 69, 287, 25, 219] and only recently was derived in acoustics [513]. Remarkably, a special case of negative recoil force, antiparallel to the radiation-pressure force, provides one of the main mechanisms of pulling forces and tractor beams in optics [118] and acoustics [166, 364] (see Section V.7).
@@
In terms of the particle’s dipole moments, the optical recoil force (24) can be maximized when the electric and magnetic dipole moments are orthogonal, in phase, and with their amplitudes related as . These are called the Kerker conditions [347, 299], whereas such an electromagnetic source is termed a Huygens source [379]. Under these conditions, the radiation far fields of the two dipole moments interfere constructively in the direction of and destructively in the opposite direction, Fig. 8(b). Most commonly, the Rayleigh-particle polarizabilities can be designed such that, in the incident plane-wave field with a wave-vector , the Kerker condition is achieved with a maximized forward scattering, [299], thereby providing a backward pulling force [143], or with maximum back-scattering, (the anti-Kerker condition) [18, 226].
@@
An acoustic analogue of the Kerker condition is provided by the acoustic monopole and dipole moments being in phase and with amplitudes [567, 353, 582]. Their radiation diagrams interfere constructively in the direction of , and destructively in the opposite direction, Fig. 8(c), enhancing the corresponding acoustic recoil force. Similar destructive interference can also happen for higher order multipoles, often referred to as the generalized Kerker conditions (see Section IV.6).
@@
In terms of the optical incident-field properties, the real and imaginary parts of the complex Poynting momentum density in Eq. (24) are generally independent from the canonical momentum densities and the energy density gradients in structured fields. Due to this, the recoil force can have a nonzero component in the ‘third’ direction, orthogonal both to the radiation and gradient forces (22) [87, 69, 25, 345, 492, 592, 623]. In particular, using the Belinfante-Rosenfeld relation (1) for , one can see that this quantity differs from the canonical momentum density by a curl of the spin density. Therefore, in contrast to the artificially-introduced ‘spin-curl’ term in the radiation-pressure force [13], the recoil force does have a polarization-dependent spin-curl component [87, 69, 25, 345, 22, 23, 492].
@@
In acoustics, using the relation , we note that the recoil force (24) is determined by the quantities and . Therefore, these recoil force terms are always aligned with the corresponding gradient and radiation-pressure terms in Eqs. (22), and cannot produce a force in the ‘third’ direction.
@@
Thus, the recoil torques are also associated with the corresponding spin densities in the incident wave field. Furthermore, using Eqs. (III.3), one can see that the proportionality coefficients are determined by the corresponding scattering cross-sections: . Adding the recoil torques (25) to Eqs. (23), where the proportonality coefficients are , and using , we obtain the total torque as
These important results state that the optical and acoustic torques on isotropic particles are determined by the corresponding spin densities and the absorption cross-sections [365, 401, 400, 543, 513]. Therefore, these torques vanish identically for isotropic lossless particles, such as dielectric particles in an optical field, and this result stays true for arbitrary-size particles [365, 614]. This is in contrast to the radiation-pressure forces in Eqs. (22), which are determined by the extinction cross-sections, and do not vanish even for lossless isotropic particles.
Finally, adding the recoil terms of the absorption rate from Table 2 to the ‘mixed terms’ absorption rates calculated in (21), we find that the total absorption rates are also determined by the absorption cross-sections:
There is a seeming paradox related to these results. According to the general Eqs. (12), the absorption rate characterizes the energy transfer from field to particle. For a lossless isotropic particle, the absorption rates (27) vanish. However, the radiation-pressure force in Eq. (22) (proportional to the extinction cross-section ), as well as the gradient and recoil forces in Eqs. (22) and (24) do not vanish for a lossless isotropic particle. Thus, these radiation forces can accelerate the particle, thereby increasing its kinetic energy. The resolution of this paradox lies in the general assumption that we deal with purely monochromatic incident and scattered fields, and hence neglect the Doppler effect of a finite velocity of the particle on the scattered field. As a result, Eqs. (27) rather describe the thermal energy absorbed by the particle, while changes in the particle’s kinetic energy require taking into account non-monochromaticity of the scattered wavefield.
@@ IV Complex forces and torques
We are now in the position to consider more sophisticated optical and acoustic interactions with particles, beyond the isotropic Rayleigh case.
@@ IV.1 Anisotropic particles
We first consider anisotropic Rayleigh particles. Anisotropy can originate either from its non-spherical shape or from the anisotropy of the material (described by tensor parameters , , , ). In any case, if the particle is small, , its interaction with the wave field can still be described by the dipole and monopole moments, but the dipole polarizabilities (III.3) now become tensors, which describe all the anisotropic properties of the particle:
@@
This means that the dipole moments induced by the incident wave field are generally not aligned with this field. Importantly, all equations in Table 2 remain valid, but with the tensor polarizabilities these do not result in Eqs. (21)–(23) and hence cannot be simply expressed via basic dynamical properties of the incident field. Thus, for anisotropic particles, the gradient forces, radiation forces, and torques are not generally aligned with the the gradient of the energy, canonical momentum, and spin densities.
@@
Since the generic dipole-polarizability tensors have many degrees of freedom, one has to consider each specific case of an anisotropic particle in a given incident field. Even the simplest uniaxial-type anisotropy (such as spheroid particles) in the simplest plane-wave incident field can produce qualitatively new types of forces and torques with new functionalities [513, 392].
@@
Some general remarks can also be made for the torques on anisotropic particles. Recall that transmission of a polarized electromagnetic wave through an anisotropic plate underlay pioneering works on the optical spin angular momentum and torque [468, 433, 80, 81, 265, 111]. The main effect of an anisotropic material is that it changes polarization of the scattered (e.g., transmitted) light. Since the spin angular momentum of the field (5) or (10) is determined by the degree of its circular polarization (Fig. 4), changes in the polarization of the scattered light affect the angular momentum balance and hence generate torques. Thus, in contrast to the torques on isotropic particles, Eqs. (23), (25), and (26), which occur due to absorption of the spin angular momentum, the torque on anisotropic particle can exist without absorption, from changes in the scattered-field spin. Within this mechanism, one can distinguish two cases, described already in [468] and observed in [213].
@@
First, a circularly- or elliptically-polarized wave interacting with an anisotropic particle can produce a constant torque, similar to the rotating half-wave plate in experiments [80, 265, 81, 111]. Assuming a non-absorbing particle with a diagonal dipole-polarizability tensor and the corresponding incident-field spin aligned with the -axis, this torque has the form [542]
@@
This equation is derived from the torque formulas in Table 2 supplied with Eqs. (III.3) for a non-absorbing particle (see Appendix E).
@@
Second, for a linearly- or elliptically-polarized incident wave, there can be a torque which vanishes at certain orientation of the particle. Such torque acts to align the particle with the polarization of the incident field [213, 513]. This torque appears even for zero spin of the incident wave, because the scattered/transmitted field has a non-zero spin when the incident polarization is not aligned with the optical axes of the plate. For a diagonal polarizability tensor and a linear incident-field polarization tilted with respect to the -axis by an angle within the -plane, such torque has the form [513, 452]
This equation is derived straightforwardly from the main torque terms in Table 2.
@@
Remarkably, although these anisotropy-induced torques have been considered exclusively in optics (apart from [513]), the above considerations are universal for any types of vector fields and dipoles: electric, magnetic, or acoustic. Equations (29) and (30) involve the corresponding dipole polarizability tensors and the incident-field polarization.
@@
Akin to the torques originating from the change in the spin angular momentum of light transmitted through an anisotropic particle/plate, there are also torques on more complicated particles/plates generating orbital angular momentum (OAM) in the transmitted or reflected wave. Such wave elements have either space-variant anisotropy axes (q-plates) or anisotropic chiral shapes (spiral-phase plates). The corresponding optical and acoustic torques on these elements were observed in [253, 587].

### v3

@@ III.3 Isotropic Rayleigh particles
Here, the complex-valued quantities are the corresponding polarizabilities of the particle, which are defined such that all have the dimension of volume: . We write relations (III.3) in the general form using the polarizability matrix, which is diagonal in this case. Since there are different conventions for defining electromagnetic and acoustic dipole and monopole moments, we present the fields produced by the moments adopted in this review in Appendix A.
The polarizabilities can be connected with the scattering, absorption, and extinction cross-sections, related as . In particular, for all polarizabilities in Eqs. (III.3), we have:
@@
These relations follow from the direct calculation of the ‘mixed’ and ‘scattered’ energy fluxes through a closed surface around the particle, normalized to the absolute value of the ‘incident’ energy flux density, assuming an incident plane-wave field [607, 341, 602, 112]. For particles without gain, and , where equality holds in the lossless case. These relations impose constraints on the polarizabilities, sometimes referred to as the optical theorem [444, 75, 476]. Note also that the absorption cross-section equals the normalized absorption rate: , in the case of an incident plane wave.
@@
Explicit formulas for the isotropic Rayleigh-particle polarizabilities in terms of the material parameters can be derived from the electromagnetic Mie theory and its acoustic analogue. The results are summarized in Appendix B and Table 5.
@@
Substituting Eqs. (III.3) into the ‘mixed’ terms (linear in the dipole and monopole moments) in Table 2, we obtain the absorption rates, forces, and torques on the particle conveniently expressed via the energy, canonical momentum, and spin densities of the incident wave field, Eqs. (2), (6), (7), (9), (11), (12), and (13) (summarized in Table 1):
@@
These are the most important optical and acoustic forces acting on small particles, schematically illustrated in Fig. 7. The remarkable symmetry between optical and acoustic quantities justifies the unified framework employed here. In Eqs. (24), one can distinguish between the gradient forces , determined by gradients of the corresponding energy-density terms, and the radiation-pressure forces , governed by the canonical momentum density terms. Notably, in paraxial optical fields, the energy density gradients are nearly orthogonal to the canonical momentum (propagation) direction, causing the gradient and radiation-pressure forces to act in nearly perpendicular directions. The torques (25) are naturally determined by the corresponding spin densities. Importantly, Eqs. (23)–(25) are valid for an arbitrarily inhomogeneous (structured) incident field, including near-fields and evanescent waves.
@@
The forces and torques (24) and (25) underpin the optical and acoustic manipulation of small particles, enabling trapping, tweezers, spanners, levitation, and related techniques. [190, 423, 261, 114, 41, 639, 187, 578, 232, 672, 619, 247, 460, 467, 394, 675, 663, 123, 410, 450, 593, 65, 370, 224, 350, 193, 286, 646, 188, 339, 254, 127, 684, 541] Pioneering studies of the dipole radiation forces were performed by Gordon and Ashkin in 1980 [257, 47] (considering interaction of an electric-dipole particle with an incident electromagnetic wave) and by Gor’kov in 1961 for the acoustic gradient force [258]. Since then, numerous works have extended these results [144, 439, 511, 190, 578, 73, 91, 105, 123, 602, 554, 540], until they acquired the modern form (24). Surprisingly, Eqs. (25) for optical and acoustic torques were derived only relatively recently [145, 133, 91, 105, 438, 554, 602].
@@
At lowest order, the imaginary parts of the polarizabilities in Eqs. (23)–(25) are commonly associated with absorption in the particle. However, according to Eqs. (III.3), they are in fact determined by the extinction cross-sections: which include both absorption and scattering. This results in nonzero imaginary parts of the polarizabilities even for lossless particles with : , Eqs. (III.3) (see Appendix B). In particular, this means that the radiation-pressure forces in Eqs. (24) do not vanish for such lossless particles [269]. This phenomenon is sometimes termed “radiation friction” [562, 439, 561, 13, 341, 602]. In contrast, the absorption rates and torques on lossless isotropic particles do vanish, but demonstrating this rigorously requires taking into account the contributions from the ‘scattered’ fluxes, which are analyzed in the next Section III.4.
@@
It is worth noting that some studies of optical forces [14, 394, 305, 232, 625, 439, 578] use an alternative yet equivalent form of the electromagnetic radiation-pressure force in Eq. (24). This form is obtained by expressing the canonical momentum densities via the difference of the kinetic Poynting momentum (3) and curls of the corresponding spin densities, according to the Belinfante-Rosenfeld relation (1): . However, this decomposition, meant to introduce a term proportional to the Poynting vector and an additional “spin-curl” force, can be delusive. Indeed, the canonical momentum density has a clear physical interpretation as a natural measure of the mean local phase gradient (weighted by the field intensity) [79, 81, 91, 105], universal for any wave field [108] and independent of polarization in paraxial uniformly-polarized fields. In contrast, decomposing the force into Poynting-vector and spin-curl contributions introduces two polarization-dependent force terms even in simple paraxial fields, which partially cancel each other, resulting ultimately in the same radiation-pressure force.
@@ III.4 Recoil force and torque
We now consider the second-order forces and torques arising from the pure scattering-field fluxes, which are quadratic in the dipole and monopole moments. These contributions are labeled in Table 2 as “scattered (recoil)”. Figure 8 schematically illustrates the nature of the recoil force. It originates from the asymmetry of the scattered momentum flux, which in turn stems from the interference of the symmetric electric and magnetic dipole or acoustic dipole and monopole contributions [439, 567, 627] (see also Appendix D).
@@
where and are the natural complex extensions of the electromagnetic and acoustic kinetic momentum densities (3) and (10).
@@
The recoil force (26) has been studied in optics since 2009 [145, 439, 73, 91, 72, 304, 27, 232], and only recently it was derived in acoustics [567]. Remarkably, a special case of negative recoil force, antiparallel to the radiation-pressure force, provides one of the main mechanisms of pulling forces and tractor beams in optics [124] and acoustics [175, 396] (see Section V.7).
@@
In terms of the particle’s dipole moments, the optical recoil force (26) is maximized when the electric and magnetic dipole moments are orthogonal, in phase, and have amplitudes satisfying . These relations are known as the Kerker conditions [377, 318], and such an electromagnetic source is termed a Huygens source [414]. Under these conditions, the far-field radiation patterns of the two dipoles interfere constructively in the direction of and destructively in the opposite direction, Fig. 8(b). Most commonly, the Rayleigh-particle polarizabilities can be designed such that, in an incident plane-wave field with a wave-vector , the Kerker condition is achieved with maximal forward scattering, [318], thereby producing a backward pulling force [151]. Conversely, the anti-Kerker condition, , yields maximal back-scattering [20, 239].
@@
An acoustic analogue of the Kerker condition occurs when the acoustic monopole and dipole moments are in phase and have amplitudes satisfying [627, 384, 642]. Their radiation diagrams interfere constructively in the direction of , and destructively in the opposite direction, Fig. 8(c), enhancing the acoustic recoil force. Similar interference effects can also happen for higher-order multipoles, often referred to as the generalized Kerker conditions (see Section IV.6).
@@
From the perspective of the incident optical field, the real and imaginary parts of the complex Poynting momentum density in Eq. (26) are generally independent of the canonical momentum densities and the energy density gradients in structured fields. Therefore, the recoil force can have a nonzero component in a ‘third’ direction, orthogonal to both the radiation and gradient forces (24) [91, 72, 27, 375, 545, 654, 691, 149, 542]. In particular, using the Belinfante-Rosenfeld relation (1) for , one can see that this quantity differs from the canonical momentum density by a curl of the spin density. Thus, in contrast to the artificially-introduced ‘spin-curl’ term in the radiation-pressure force [14], the recoil force does contain a polarization-dependent spin-curl component [91, 72, 27, 375, 24, 25, 545]. Notably, a transverse spin-dependent recoil force can also arise in a simple plane-wave field interacting with a specially designed anisotropic object side-scattering different spin components in opposite directions [389].
@@
In acoustics, using the relation , the recoil force (26) is determined by the quantities and . Therefore, these recoil-force terms are always aligned with the corresponding gradient and radiation-pressure terms in Eqs. (24), and cannot produce a force in the ‘third’ direction.
@@
Thus, the recoil torques are also associated with the corresponding spin densities of the incident wave field. Furthermore, from Eqs. (III.3), one can see that the proportionality coefficients are determined by the corresponding scattering cross-sections: . Adding the recoil torques (27) to Eqs. (25), where the proportonality coefficients are , and using , we find the total torque:
These important results show that the optical and acoustic torques on isotropic particles are determined by the corresponding spin densities and the absorption cross-sections [397, 438, 437, 602, 567]. Consequently, the torque vanishes identically for isotropic lossless particles, such as dielectric particles in an optical field, and this remains true for particles of arbitrary size [397, 679]. This contrasts with the radiation-pressure forces in Eqs. (24), which are proportional to the extinction cross-sections and do not vanish even for lossless isotropic particles.
Finally, adding the recoil terms in the absorption rates from Table 2 to the ‘mixed terms’ (23), we find that the total absorption rates are also determined by the absorption cross-sections:
A seeming paradox arises from these expressions. According to the general Eqs. (14), the absorption rate quantifies the energy transfer from the field to the particle. For a lossless isotropic particle, the absorption rates (29) vanish. However, the radiation-pressure force in Eq. (24) (proportional to the extinction cross-section ), as well as the gradient and recoil forces in Eqs. (24) and (26) do not vanish in this case. These forces can accelerate the particle and increase its kinetic energy. The resolution of this paradox is that the above derivations assume perfectly monochromatic incident and scattered fields, neglecting Doppler effects due to the finite velocity of the particle. Under this assumption, Eqs. (29) describe only the thermal energy absorbed by the particle, whereas changes in its kinetic energy require accounting for non-monochromaticity of the scattered field.
@@ IV Complex Forces and Torques
We are now in a position to consider more sophisticated optical and acoustic interactions with particles, beyond the isotropic Rayleigh case.
@@ IV.1 Anisotropic particles
We first examine anisotropic Rayleigh particles. Anisotropy can originate either from a non-spherical shape of the particle or from the intrinsic anisotropy of its material (described by tensor parameters , , , and ). When the particle is small, , its interaction with the wave field can still be described in terms of the dipole and monopole moments, but the dipole polarizabilities (III.3) now become tensors, which capture the anisotropic response of the particle:
@@
As a consequence, the induced dipole moments are generally not aligned with the local incident fields. Importantly, all equations in Table 2 remain valid, but with the tensor polarizabilities they no longer reduce to Eqs. (23)–(25). Hence, the resulting forces and torques cannot be expressed via the basic dynamical properties of the incident field. Thus, for anisotropic particles, the gradient forces, radiation forces, and torques are generally not aligned with the energy-density gradients, canonical momenta, and spin densities.
@@
Since generic dipole-polarizability tensors carry many degrees of freedom, one has to analyze each case of an anisotropic particle in a given incident field individually. Even the simplest uniaxial anisotropy (e.g., spheroidal particles) in a uniform plane wave can produce qualitatively new types of forces and torques, enabling new functionalities [567, 429].
@@
Some general remarks can also be made regarding torques on anisotropic particles. Recall that the transmission of a polarized electromagnetic wave through an anisotropic plate underpinned the pioneering studies of optical spin angular momentum and torque [516, 472, 83, 84, 280, 116] (see Section I.1). The main effect of an anisotropic material is that it modifies the polarization of the scattered (e.g., transmitted) light. Since the spin angular momentum of the field (7) or (12) is determined by the degree of circular polarization (Fig. 4), changes in the polarization of the scattered field affect the angular momentum balance and thereby generate torques. In contrast to the case of isotropic particles, where torques (25), (27), and (28) arise only from absorption of the spin angular momentum, torques on anisotropic particle can exist without absorption, due to changes in the scattered-field spin. Within this mechanism, two distinct situations can be identified, already described in [516] and experimentally observed in [226].
@@
First, when a circularly or elliptically polarized wave interacts with an anisotropic particle, a constant torque can be generated, similar to the torque on a half-wave plate in experiments [83, 280, 84, 116] or on an electric-dipole antenna in [18]. For a non-absorbing particle with a diagonal dipole-polarizability tensor and incident-field spin aligned with the -axis, this torque has the form [597]
@@
This equation follows from the torque formulas in Table 2 supplied with Eqs. (III.3) for a non-absorbing particle (see Appendix E).
@@
Second, for a linearly or elliptically polarized incident wave, there is a torque that vanishes at certain orientations of the particle. Such torque acts to align the particle with the polarization of the incident field [226, 567]. This torque appears even when the incident spin vanishes, because the scattered/transmitted field acquires a non-zero spin when the incident polarization is not aligned with the particle’s optical axes. For a diagonal polarizability tensor and a linear polarization of the incident field, tilted by an angle with respect to the -axis in the -plane, the torque is given by [567, 495]
This result follows directly from the main torque contributions in Table 2.
@@
Remarkably, although these anisotropy-induced torques have been considered exclusively in optics (apart from [567]), the above considerations are universal for any type of vector field and associated dipoles: electric, magnetic, or acoustic. Equations (31) and (32) involve the corresponding dipole polarizability tensors and the incident-field polarization.
@@
Similarly to torques originating from the change in the spin angular momentum of light transmitted through an anisotropic particle/plate, there are also torques on more complex particles/plates that generate orbital angular momentum (OAM) in the transmitted or reflected wave. Such wave elements have either space-variant anisotropy axes (q-plates) or anisotropic chiral geometries (spiral-phase plates). The corresponding optical and acoustic torques on such structures have been observed in [267, 649, 389].